Operations Report – February 15th

Crew 205 Operations Report 15 – FEB – 2019

SOL: 5

Name of person filing report: Natalia Larrea

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned off at 0830 (15th Feb),Turned on at 1730 (15th Feb),

Solar— SOC % – Turned off Gen (15th Feb) 69%, Turned on Gen (15th Feb) 74%

Diesel Reading – 97%

Propane Reading – ~ 45%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 64.5%, ~ 355 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 240.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01412441 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable566

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 50.1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 74.3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 73.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: Yes, number 2

Reason for use: exploration

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: 40 min

Notes on ATVs: None

HabCar used and why, where? Not used

Crew Car used and why, where? Not used

General notes and comments: Downstairs furnace was turned off due to higher ambient temperatures.

Summary of internet: Nothing to report

Summary of suits and radios: Two radios have been stuck in privacy and NOAA weather radio mode, respectively since we arrived. Following the radio’s manual did not fix them. Notes describing the issues have been clipped to the radios to assist future troubleshooting and identify them as not working. We still have plenty of radios working so it is not urgent.

Summary of Hab operations: Replaced the air filter in the heater in the lower-deck. Heater upstairs is nominal and currently used.

Summary of Green Hab operations: GreenHab smoke detector was beeping (it seems it was low in battery). Replacing its AA batteries fixed the problem.

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: We observed a water leak above the workbench (we found a water puddle this morning)

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Due to strong winds a white canvas flew towards the tunnels. We believe it belongs to Director’s compound.

Copyright © The Mars Society. All rights reserved. | Main Site